OpenStudy (anonymous):

write an equation in point slope form of the line that passes through the point (0,b) and have the slope of m. explain how slope intercept form is a particular case of point slope form?

OpenStudy (anonymous):

Well the equation is simply y=mx+b

OpenStudy (anonymous):

Because:

OpenStudy (anonymous):

(y-b)=m(x+0) y-b=mx+0 +b +b y=mx+b

OpenStudy (anonymous):

But i don't know how to explain how slope-intercept is a particular case of point-slope form.... I hope it helped somewhat

OpenStudy (anonymous):

Do you get what I did to solve the first problem?